OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2π). cosx=sin2x

OpenStudy (anonymous):

cosx=0 x=(2n+1)pia/2 2sinx-1=o sinx=1/2 x=pia/6=30